Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below:

Sales $ 18,100,000
Net operating income $ 5,900,000
Average operating assets $ 35,400,000

Required:

1. Compute the margin for Alyeska Services Company. (Round your answer to 2 decimal places.)

2. Compute the turnover for Alyeska Services Company. (Round your answer to 2 decimal places.)

3. Compute the return on investment (ROI) for Alyeska Services Company. (Round your intermediate calculations and final answer to 2 decimal places.)
Answer #1

Margin = Net operating income / sales

= 5,900,000/18,100,000

= 32.60%

Turnover = Sales / Average operating assets

=18,100,000/35,400,000

= 0.51

Roi = Margin * turnover

= 32.6*0.51

= 16.63%

Comment if you face any issues